OpenStudy (anonymous):

Find how many arrangements there are possible from the word STUDIO a. The letter S is not the first letter b. The letter O is not the first or last letter

terenzreignz (terenzreignz):

\[\Large \left[\begin{matrix}S &T&U&D&I&O\\?&?&?&?&?&?\end{matrix}\right]\]

terenzreignz (terenzreignz):

You can think of it as counting the number of 'slots' each letter may go to. \[\Large \left[\begin{matrix}S &T&U&D&I&O\\1&2&3&4&5&6\end{matrix}\right]\]

OpenStudy (anonymous):

so 5!*6!?

terenzreignz (terenzreignz):

That's way too much... Even more than the number of rearrangements of STUDIO without restrictions... Just hear me out first :P

OpenStudy (anonymous):

sure

terenzreignz (terenzreignz):

Now, let's consider the first restriction, namely, that S does not take the first position, or in other words, it may not take position#1 In that case, how many 'slots' are available for S?

OpenStudy (anonymous):

5

terenzreignz (terenzreignz):

That's right :) Now, once S is placed, how many 'slots' are available for T? There is no restriction for T, however, one of the slots is now currently occupied by S...

OpenStudy (anonymous):

5

terenzreignz (terenzreignz):

That makes 5*5 possibilities so far... Catch me so far? What about the letter U?

OpenStudy (anonymous):

4

terenzreignz (terenzreignz):

And it goes on, right? :) 5*5*4*3*2*1 or \[\Large 5\cdot 5!\]

OpenStudy (anonymous):

ok i got that

terenzreignz (terenzreignz):

I'm sorry, there IS an easier way to do that... First you count the number of rearrangements without restrictions, aye? that's exactly 6!

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

And then, you count precisely those rearrangements that DO have S on the first position, and that's like placing S on the first position and arranging the rest (there are 5 remaining) and that's 5!

OpenStudy (anonymous):

?

terenzreignz (terenzreignz):

We count the ones that don't obey the restriction. Namely, the number of rearrangements that DO have S on the first position. So that means, we fix S on the first position... \[\Large \left[\begin{matrix}S &T&U&D&I&O\\S&?&?&?&?&?\end{matrix}\right]\]

terenzreignz (terenzreignz):

And then rearrange the remaining letters on the remaining five slots, and that gives 5! ways, right?

terenzreignz (terenzreignz):

I'll stop if you find this more confusing ^_^ Just let me know...

terenzreignz (terenzreignz):

\[\Large 5\cdot 5!\]IS the answer for the first bit, I'm just showing you another way to get it.

OpenStudy (anonymous):

i think i understood the first method better.

terenzreignz (terenzreignz):

Yeah...it is more down-to-earth. Oh well :D Let's do the second?

OpenStudy (anonymous):

For the next part i came up with 4.4! but its not the correct answer. i used the first method

terenzreignz (terenzreignz):

Well, you can't just do these things methodically... :D

terenzreignz (terenzreignz):

Here... \[\Large \left[\begin{matrix}S &T&U&D&I&O\\1&2&3&4&5&6\end{matrix}\right]\]

OpenStudy (anonymous):

there are a total of 4 positions for O

terenzreignz (terenzreignz):

That's right. So once O is placed, let's place the rest :D

terenzreignz (terenzreignz):

But then again, the next letter will STILL have 5 possibilities for it, since only the letter O has any restrictions at all...

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

Right? ^_^ That was your error.

OpenStudy (anonymous):

so 4.5!?

terenzreignz (terenzreignz):

mhmm :) \[\Large 4\cdot 5!\]

OpenStudy (anonymous):

i think my main concern is not understanding the concept well...

terenzreignz (terenzreignz):

Combinatorics calls for concepts a lot. Practice, practice, practice! :D

OpenStudy (anonymous):

oh i've heard of that before eh?

OpenStudy (anonymous):

thanks

terenzreignz (terenzreignz):

Only naturally :D